Two electrons in an atom are in 1s (n=1, l = 0) and 2p (n=1, l = 1) states. What are the possible values of the total angular momentum quantum number for the individual particles j1 and j2? What are the possible values for jT, the quantum number for the sum of the two total angular momenta (jT = j1 + j2).
Answer #1

Total angular momentum quantum number have the value between

| l-s| <= j <= | l+s |

For 1s, l= 0, s= 1/2

j1 = 1/2 , 1/2

For 2p, l= 1, s= 1/2

j= 1/2, 3/2